Lecture Notes in Computer Science: Authors’ Instructions for the Preparation of Camera-Ready Contributions to LNCS/LNAI Proceedings

Alfred Hofmann^{1}, Antje Endemann^{1}, Ingrid Beyer^{1}, Karin Henzold^{1}, Anna Kramer^{1}, Erika Siebert-Cole^{1}, Angelika Bernauer-Budiman^{2}, Martina Wiese^{2}, and Anita Bürk^{3}

1 Springer-Verlag, Computer Science Editorial III, Postfach 10 52 80, 69042 Heidelberg, Germany

 {Hofmann, Endemann, Beyer, Henzold, Kramer, Erika.Siebert-Cole,
                           LNCS}@Springer.de
            http://www.springer.de/comp/lncs/index.html

2 Springer-Verlag, Computer Science Production, Postfach 10 52 80, 69042 Heidelberg, Germany

                    {Bernauer, Wiese}@Springer.de

3 Springer-Verlag, Marketing Management, Postfach 10 52 80, 69042 Heidelberg, Germany

                           Buerk@Springer.de

Abstract. The abstract should summarize the contents of the paper and should contain at least 70 and at most 150 words. It should be set in 9-point font size and should be inset 1.0 cm from the right and left margins. There should be two blank (10-point) lines before and after the abstract. …

1 Introduction

The preparation of manuscripts which are to be reproduced by photo-offset requires special care. Papers submitted in a technically unsuitable form will be returned for retyping, or canceled if the volume cannot otherwise be finished on time.

1.1 LNCS Online

Springer-Verlag now provides the full-text version of the LNCS and LNAI proceedings online. Therefore please submit to the volume editors (and not to Springer-Verlag), together with your own single-sided printout of the final version of your contribution (which cannot be modified at a later stage), your source (input) files, e.g. TEX files for the text and PS or EPS files for figures, the final DVI file (for papers prepared using L^{A}TEX or TEX), the final PS file^{4}, and, if possible, a PDF file of the final version of your contribution. If you have prepared your paper using a text processing system other than L^{A}TEX or TEX, please also submit RTF files. Make sure that the text is identical in all cases.

4 When generating the PS file please avoid using the option “reverse order”.

2 Manuscript Preparation

You are strongly encouraged to use L^{A}TEX2_{ε} for the preparation of your camera-ready manuscript together with the corresponding Springer class file llncs.cls; see Sect. 3. Only if you use L^{A}TEX2_{ε} can hyperlinks be generated in the online version of your manuscript.

If you are unable to use L^{A}TEX, you may use MS Word together with the template svlncs.dot (see Sect. 4) or any other text processing system. In the latter case, please follow these instructions closely in order to make the volume look as uniform as possible.

We would like to stress that the class/style files and the template should not be manipulated and that the guidelines regarding font sizes and format should be adhered to. This is to ensure that the end product is as homogeneous as possible.

2.1 Printing Area

The printing area is 122 mm × 193 mm. The text should be justified to occupy the full line width, so that the right margin is not ragged, with words hyphenated as appropriate. Please fill pages so that the length of the text is no less than 180 mm.

2.2 Layout, Typeface, Font Sizes, and Numbering

Use 10-point type for the name(s) of the author(s) and 9-point type for the address(es) and the abstract. For the main text, please use 10-point type and single-line spacing. We recommend using Computer Modern Roman (CM) fonts, Times, or one of the similar typefaces widely used in photo-typesetting. (In these typefaces the letters have serifs, i.e., short endstrokes at the head and the foot of letters.) Italic type may be used to emphasize words in running text. Bold type and underlining should be avoided. With these sizes, the interline distance should be set so that some 45 lines occur on a full-text page.

Headings. Headings should be capitalized (i.e., nouns, verbs, and all other words except articles, prepositions, and conjunctions should be set with an initial capital) and should, with the exception of the title, be aligned to the left. Words joined by a hyphen are subject to a special rule. If the first word can stand alone, the second word should be capitalized. The font sizes are given in Table 1.

Here are some examples of headings: “Criteria to Disprove Context-Freeness of Collage Languages”, “On Correcting the Intrusion of Tracing Non-deterministic Programs by Software”, “A User-Friendly and Extendable Data Distribution System”, “Multiflip Networks: Parallelizing GenSAT”, “Self-determinations of Man”.

Lemmas, Propositions, and Theorems. The numbers accorded to lemmas, propositions, and theorems etc. should appear in consecutive order, starting with the number 1, and not, for example, with the number 11.

Table 1. Font sizes of headings. Table captions should always be positioned above the tables. The final sentence of a table caption should end without a period

Heading level  Example  Font size and style
Title (centered) Lecture Notes … 14 point, bold
1st-level heading 1 Introduction  12 point, bold
2nd-level heading 2.1 Printing Area  10 point, bold
3rd-level heading Headings. Text follows …  10 point, bold
4th-level heading Remark. Text follows …  10 point, italic

2.3 Figures and Photographs

Please produce your figures electronically, if possible, and integrate them into your text file. For L^{A}TEX users we recommend using the style files psfig or epsf (see Sect. 3).

Check that in line drawings, lines are not interrupted and have constant width. Grids and details within the figures must be clearly readable and may not be written one on top of the other. Line drawings should have a resolution of at least 800 dpi (preferably 1200 dpi). For digital halftones 300 dpi is usually sufficient. The lettering in figures should have a height of 2 mm (10-point type). Figures should be scaled up or down accordingly. Please do not use any absolute coordinates in figures. If possible, the files of figures (e.g. PS files) should not contain binary data, but be saved in ASCII format.

If you cannot provide your figures electronically, paste originals into the manuscript and center them between the margins. For halftone figures (photos), please forward high-contrast glossy prints and mark the space in the text as well as the back of the photos clearly, so that there can be no doubt about where or which way up they should be placed.

Figures should be numbered and should have a caption which should always be positioned under the figures, in contrast to the caption belonging to a table, which should always appear above the table. Please center the captions between the margins and set them in 9-point type (Fig. 1 shows an example). The distance between text and figure should be about 8 mm, the distance between figure and caption about 5 mm.

If possible (e.g. if you use L^{A}TEX) please define figures as floating objects. L^{A}TEX users, please avoid using the location parameter “h” for “here”. If you have to insert a pagebreak before a figure, please ensure that the previous page is completely filled.

Remark 1. In the printed volumes, illustrations are generally black and white (halftones), and only in exceptional cases, and if the author is prepared to cover the extra cost for color reproduction, are color pictures accepted. If color illustrations are necessary, please send us color-separated files if possible. Color pictures are welcome in the electronic version at no additional cost.

Remark 2. To ensure that the reproduction of your illustrations is of reasonable quality we advise against the use of shading. The contrast should be as pronounced as possible. This particularly applies for screenshots.

Fig. 1. One kernel at x_{s} (dotted kernel) or two kernels at x_{i} and x_{j} (left and right) lead to the same summed estimate at x_{s}. This shows a figure consisting of different types of lines. Elements of the figure described in the caption should be set in italics, in parentheses, as shown in this sample caption. The last sentence of a figure caption should generally end without a period

2.4 Formulas

Displayed equations or formulas are centered and set on a separate line (with an extra line or halfline space above and below). Displayed expressions should be numbered for reference. The numbers should be consecutive within each section or within the contribution, with numbers enclosed in parentheses and set on the right margin. For example,

ψ(u) = ∫_{o}^{T} [\frac{1}{2} (Λ_{o}^{−1}u, u) + N^{∗}(−u)] dt .  (1)

Please punctuate a displayed equation in the same way as ordinary text but with a small space before the end punctuation. L^{A}TEX users can find more examples of how to typeset equations in the file llncs.dem (see Sect. 3).

2.5 Program Code

Program listings or program commands in the text are normally set in typewriter font, e.g., CMTT10 or Courier.

Example of a Computer Program

program Inflation (Output)
  {Assuming annual inflation rates of 7%, 8%, and 10%,...
   years};
   const
      MaxYears = 10;

   var
      Year: 0..MaxYears;
      Factor1, Factor2, Factor3: Real;
   begin
      Year := 0;
      Factor1 := 1.0; Factor2 := 1.0; Factor3 := 1.0;
      WriteLn('Year  7% 8% 10%'); WriteLn;
      repeat
        Year := Year + 1;
        Factor1 := Factor1 * 1.07;
        Factor2 := Factor2 * 1.08;
        Factor3 := Factor3 * 1.10;
        WriteLn(Year:5,Factor1:7:3,Factor2:7:3,Factor3:7:3)
      until Year = MaxYears
end.

(Example from Jensen K., Wirth N. (1991) Pascal user manual and report. Springer, New York)

2.6 Footnotes

The superscript numeral used to refer to a footnote appears in the text either directly after the word to be discussed or – in relation to a phrase or a sentence – following the punctuation sign (comma, semicolon, or period). Footnotes should appear at the bottom of the normal text area, with a line of about 2 cm in TEX and about 5 cm in Word set immediately above them.^{5}

2.7 Citations

The list of references is headed “References” and is not assigned a number in the decimal system of headings. The list should be set in small print and placed at the end of your contribution, in front of the appendix, if one exists. Please do not insert a pagebreak before the list of references if the page is not completely filled. An example is given at the end of this information sheet. For citations in the text please use square brackets and consecutive numbers: [1], [2], [3] …

2.8 Page Numbering and Running Heads

Your paper should show no printed page numbers; these are allocated by the volume editor. Please indicate the ordering of your pages by numbering the sheets in pencil at the bottom of the reverse side. Do not set running heads.

5 The footnote numeral is set flush left and the text follows with the usual word spacing. Second and subsequent lines are indented. Footnotes should end with a period.

2.9 Printing Quality

For reproduction we need sheets which are printed on one side only. Please use a highresolution printer, preferably a laser printer with at least 300 dpi. We prefer the text to be centered on the pages (i.e., equal margins left and right and top and bottom). The format of the paper (A4, Letter, etc.) is irrelevant.

3 Using L^{A}TEX or TEX

You will get the best results and your files will be easiest to handle if you use L^{A}TEX2_{ε} for the preparation of your camera-ready manuscript together with the corresponding Springer class file llncs.cls. Only if you use L^{A}TEX2_{ε} can hyperlinks be generated in the online version of your manuscript.

If you are unable to use L^{A}TEX2_{ε} you may use one of our old macro packages llncs (for L^{A}TEX) or plncs (for TEX).

3.1 How to Access the Springer L^{A}TEX2_{ε}, L^{A}TEX, and TEX Macro Packages

For users of L^{A}TEX (or TEX) Springer-Verlag provides the macro package llncs for L^{A}TEX (or plncs for TEX). The packages can be obtained by ftp/gopher or by email as follows:

Ftp: The internet address is ftp.springer.de, the user ID is ftp or anonymous. Please enter your email address as password. The files (mentioned above) can be found in /pub/tex. In the directory

      ftp://ftp.springer.de/pub/tex/latex/llncs/latex2e

you will find all files belonging to the L^{A}TEX2_{ε} package for LNCS. llncs.dem is a sample input file which you may use as a source for your own input. llncs.doc is the documentation of the class; llncs.dvi the resulting DVI file of

   llncs.doc.

Gopher: Point your client to ftp.springer.de. Mailserver: Send an email message to svserv@vax.ntp.springer.de con-

taining the line

      get /tex/latex/llncs2e.zip to get the L^{A}TEX2_{ε} style files,

get /tex/latex/llncs.zip to get the L^{A}TEX style files, or

      get /tex/plain/plncs.zip  to get the TEX style files.

Sending help to the server prompts advice on how to interact with the mail server. The style files must be unzipped and uu-decoded before use. In case of problems in getting or uu-decoding the style files please contact springer@vax.ntp.springer.de.

3.2 Further Instructions for L^{A}TEX and TEX Users

Please always cancel any superfluous definitions that are not actually used in your text. If you do not, these may conflict with the definitions of the macro package, causing changes in the structure of the text and leading to numerous mistakes in the proofs.

When you use L^{A}TEX or TEX and our macro packages, your text is typeset automatically in Computer Modern Roman (CM) fonts. Please do not change the preset fonts. If you have to use fonts other than the preset fonts, kindly submit these with your files.

Please use the commands \label and \ref for cross-references and the commands \bibitem and \cite for references to the bibliography, to enable us to create hyperlinks at these places.

For preparing your figures electronically and integrating them into your TEX file we recommend using the style files psfig or epsf. These can be downloaded from the DANTE ftp server at the locations

   ftp://ftp.dante.de/tex-archive/graphics/psfig/psfig.sty or
   ftp://ftp.dante.de/tex-archive/systems/knuth/local/lib/epsf.tex.

These styles have always worked smoothly with our macro package. For further details about figure preparation see Sect. 2.3. In general, please refrain from using the \special command.

Remember to submit the psfig or epsf files and further style files and fonts you have used together with your source files.

4 Using MS Word

We do not encourage the use of MS Word, particularly as the layout of the pages (the position of figures and paragraphs) can change from printout to printout. Having said this, we do provide the template sv-lncs.dot to help MS Word users prepare their cameraready manuscripts and to enable us to use their source files for the online version.

The template sv-lncs.dot and its documentation can be downloaded from the LNCS

Web page at http://www.springer.de/comp/lncs/authors.html.

5 Supplementary Material

If you wish to include color illustrations in the electronic version in place of or in addition to any black and white illustrations in the printed version, please provide the volume editors with the appropriate files.

If you have supplementary material, e.g., executable files, video clips, or audio recordings, on your server, simply send the volume editors a short description of the supplementary material and inform them of the URL at which it can be found. We will add the description of the supplementary material to the online version of LNCS and create a link to your server. Alternatively, if this supplementary material is not to be updated at any stage, then it can be sent directly to the volume editors, together with all the other files.

6 Copyright Form

Until now, we have always had a very liberal policy regarding copyright. However, we have now had to introduce a copyright form, which we ask contributing authors to complete and sign. (It is sufficient if one author from each contribution signs the form on behalf of all the other authors.) The copyright form is located on our Web page at

http://www.springer.de/comp/lncs/copyrigh.html. The printed form

should be completed and signed and sent on to the volume editors either by normal mail or by fax, who then send it on to us, together with the printed manuscript.

7 Checklist

When submitting your camera-ready manuscript to the volume editors, please make sure you include the following:

– a single-sided printout (not a photocopy) of the final version of your contribution (unless otherwise specified by the volume editor),

– your source (input) files, e.g. TEX files for the text and PS or EPS files for the figures,

– RTF files (see Sect. 1.1),

– any style files, templates, and special fonts you may have used,

– the final DVI file (for papers prepared using L^{A}TEX or TEX),

– the final PS file (not in reverse order),

– if possible, a PDF file of the final version of your contribution,

– the completed and signed copyright form.

If supplementary material is available, please provide the volume editors with

– a short description of the supplementary material,

– the supplementary material itself or the URL at which it can be found,

– the files of color figures for the electronic version.

References

1. van Leeuwen, J. (ed.): Computer Science Today. Recent Trends and Developments. Lecture Notes in Computer Science, Vol. 1000. Springer-Verlag, Berlin Heidelberg New York (1995)

2. Bruce, K.B., Cardelli, L., Pierce, B.C.: Comparing Object Encodings. In: Abadi, M., Ito, T. (eds.): Theoretical Aspects of Computer Software. Lecture Notes in Computer Science, Vol. 1281. Springer-Verlag, Berlin Heidelberg New York (1997) 415–438

3. Michalewicz, Z.: Genetic Algorithms + Data Structures = Evolution Programs. 3rd edn. Springer-Verlag, Berlin Heidelberg New York (1996)

4. Baldonado, M., Chang, C.-C.K., Gravano, L., Paepcke, A.: The Stanford Digital Library Metadata Architecture. Int. J. Digit. Libr. 1 (1997) 108–121

Appendix: Springer-Author Discount

All authors or editors of Springer books, in particular authors contributing to any LNCS or LNAI proceedings volume, are entitled to buy any book published by Springer-Verlag for personal use at the “Springer-author” discount of one third off the list price. Such preferential orders can only be processed through Springer directly (and not through bookstores); reference to a Springer publication has to be given with such orders. Any Springer office may be contacted, particularly those in Heidelberg and New York:

Springer Auslieferungsgesellschaft Springer-Verlag New York Inc. Haberstrasse 7 P.O. Box 2485 69126 Heidelberg Secaucus, NJ 07096-2485 Germany USA Fax: +49 6221 345-229 Fax: +1 201 348 4505 Phone: +49 6221 345-0 Phone: +1-800-SPRINGER

(+1 800 777 4643), toll-free in USA

Preferential orders can also be placed by sending an email to

   orders@springer.de or orders@springer-ny.com.

For information about shipping charges, please contact one of the above mentioned orders departments. Sales tax is required for residents of: CA, IL, MA, MO, NJ, NY, PA, TX, VA, and VT. Canadian residents please add 7% GST. Payment for the book(s) plus shipping charges can be made by giving a credit card number together with the expiration date (American Express, Eurocard/Mastercard, Discover, and Visa are accepted) or by enclosing a check (mail orders only).